OpenStudy (usukidoll):

ode refresher and yes I know it's a separable one! dy/dx=y-y^2

OpenStudy (usukidoll):

\[\frac{dy}{dx}=y-y^2\]

OpenStudy (usukidoll):

\[dy \frac{1}{y-y^2}=dx\]

OpenStudy (usukidoll):

I know the right hand side is x + C by default. but wtheck is going with the left...

OpenStudy (usukidoll):

I don't think I can split the fraction... so I tried u subsitution which is u = y-y^2 du = 1-2y

OpenStudy (usukidoll):

but then I get something screwed up and aarrrrghhhh wtheck.

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{y(1-y)} dy=\int\limits_{}^{}1 dx\] use partial fractions for left hand side

myininaya (myininaya):

right hand side is pretty easy

OpenStudy (usukidoll):

oh of course x.x that's the A/wfsdjfsdkl B/ sfjdslkfjdskl such and such

myininaya (myininaya):

\[\frac{1}{y(1-y)}=\frac{A}{y}+\frac{B}{1-y} \text{ yeah }\]

OpenStudy (usukidoll):

A/y + B/1-y stuff right? since it's factorable in the denominator... yeah now I remember you're a godsend thank you :)

OpenStudy (usukidoll):

and then solve for A and B and such and such ok...

OpenStudy (usukidoll):